# Unveiling the inconsistency of the Proca theory with non-minimal coupling to gravity

### v2

@@ 1 Introduction
The main question that we aim to answer in this paper is – Does Proca theory non-minimally coupled to gravity have strong coupling, and if yes, how strongly coupled the theory is? Based on the results in massive gauge theories, there are three probable outcomes: First, the theory could be already well-behaved in the massless limit, despite the presence of the longitudinal modes. In contrast, the second two possibilities contain perturbative series that are singular in mass, and consequently, which have a strong coupling scale. However, either only the longitudinal mode becomes strongly coupled, or the remaining modes of the theory become strongly coupled as well. From the perspective of self-interacting theory, the first case is natural. In cubic or quartic self-interactions, the transverse modes, dof which are also present in Maxwell’s electrodynamics, remain weakly coupled even after the perturbation theory for the longitudinal modes breaks down. The second outcome, in which the transverse modes, or tensor modes – modes of the metric perturbations – become strongly coupled, however, would be unphysical. It would imply that in the presence of non-minimal coupling, one could either not be able to recover electromagnetism or discover gravitational waves.
By studying the Proca theory, non-minimally coupled to gravity in a flat background, and taking into account the metric perturbations around it, we will show that the last case takes place – both tensor and longitudinal modes become strongly coupled, yielding a physically inconsistent result. Moreover, we will show that the way to resolve this ambiguity lies in the introduction of the disformal coupling to the metric perturbations simultaneously. Such disformal transformations, which modify the original metric by adding a bilinear term in the external fields or their field-strength have gained a lot of attention in cosmology, and construction of new theories of gravity [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74]. In addition, we will show that the production of runaway modes – modes with negative frequency in curved spacetime – recently reported in [26], can be cured with the same transformations.
@@ 2 The basics of linearized gravity and Proca theory
Here, the indices are raised and lowered the indices with the Minkowski metric.
Let us now first study the dof in gravity. For this, we will decompose the metric perturbations according to the spatial rotations:
@@ 3.1 $\beta$ coupling case
Once For the longitudinal modes, this means that the first term on the left-hand side of (28) is no longer vanishing. Once the scale $L_{\beta str}$ is reached, their first-order corrections should be evaluated as:
@@
Therefore, in this section, we have found that both tensor modes and longitudinal modes become strongly coupled once the scale reaches $L_{\beta str}$. However, this is an unphysical result. It would imply that no matter how small the mass of the photon is, gravitational waves would be strongly coupled if non-minimal coupling is taken into account.
@@ 3.2 $\alpha$ coupling case
In the previous section we have seen that due to the $\beta$ interaction, both tensor and the longitudinal modes become strongly coupled. Let us now study the coupling that involves only the Ricci scalar. First of all, we can notice that in this case, at cubic order, only the gravitational potentials are included in the interaction:
@@
Thus, in this order, we can safely avoid the danger of tensor modes becoming strongly coupled, in contrast to the previous case. Therefore, the only question that remains to be answered is if the longitudinal modes will enter a strong coupling regime at quartic order. For simplicity, let us set the tensor and transverse vector modes to zero. Then, the most dominant terms in the Lagrangian are given by:
@@ 4 How to cure the tensor modes?
In the previous sections, we have studied the $\alpha$ and $\beta$ models, which involved the coupling of the vector field with the Ricci scalar and the Ricci tensor respectively. In the case of the $\alpha$ coupling, we have found that only the longitudinal modes become strongly coupled, due to the quartic self-interaction that appears after one solves the constraints. In contrast, the coupling with the Ricci tensor proved to be troubling. There, we have found that both tensor and longitudinal modes become strongly coupled, which is a physically unacceptable result.
@@
Moreover, the couplings between $\phi$, $\psi$, and $S_{i}$ and the vector modes in the last line of $\mathcal{L}_{hA}$ can be neglected as well. One can show that the longitudinal mode in these relations enters in combination with the $A_{0}$ component, such that once the temporal constraint is resolved, it is multiplied by the mass squared, making such terms sub-dominant.
Therefore, the most important terms are given by:
@@
Here, we have omitted the contribution of the longitudinal modes, which is now subdominant. By taking into account that the minimal amplitude of quantum fluctuations for the tensor modes is singular in the Planck mass, it is clear that the two types of modes will become strongly coupled only at the Planck scale, before which the perturbative approach still holds.
@@ 5 Resolving the Runaway Modes
in the lowest order, while the last term represent higher order terms that involve couplings between the vector fields with the Ricci scalar or Ricci tensor. This guarantees that now $M_{S}^{2}=M_{T}^{2}$, and thus the runaway case disappears.
@@ 6 Conclusion
In this work, we have studied Proca theory, non-minimally coupled to gravity. The cosmological aspects of this theory have recently drawn a lot of attention, making it an intriguing possibility to underline the gravitational production of dark matter and a possible candidate for generating primordial magnetic fields and inflationary candidates. By studying this theory in flat space-time, with metric perturbations also taken into account, we have found a somewhat surprising result. On the one hand, the interaction of the vector field with the Ricci scalar results in the strong coupling of the longitudinal modes. This is relatively natural to expect, as one obtains self-interaction of the longitudinal modes at the quartic order. Beyond it, the vector and tensor modes remain weakly coupled, up to the Planck length, due to the amplitude of quantum fluctuations for the tensor modes. On the other hand, the coupling of the vector modes with the Ricci tensor indicates that both tensor and longitudinal modes become strongly coupled at the same Vainsthein scale. This, however, is an unphysical result – it would imply that if one allows for the non-minimal coupling in nature, the gravitational waves would quickly lose their linear propagator. Moreover, a recent study has found a new class of the negative frequency modes in curved space-time, known as the runaway modes [26]. In this work, we have shown that its origin lies precisely in the coupling, which is problematic even in the flat background.
@@ Acknowledgements
I am grateful to Misao Sasaki for illuminating discussions and comments on the draft of the paper. I would also like to thank Edward "Rocky" Kolb for communicating his work on the Proca theory, and for very useful discussions. In addition, I would like to thank Elisa G. M. Ferreira, Shinji Mukohyama, Ryo Namba, and Ippei Obata for very useful discussions. The work of A. H. was supported by the World Premier International Research Center Initiative (WPI), MEXT, Japan.

### v3

@@ 1 Introduction
The main question that we aim to answer in this paper is – Does Proca theory non-minimally coupled to gravity have strong coupling, and if yes, how strongly coupled the theory is? Based on the results in massive gauge theories, there are three probable outcomes: First, the theory could be already well-behaved in the massless limit, despite the presence of the longitudinal modes. In contrast, the second two possibilities contain perturbative series that are singular in mass, and consequently, which have a strong coupling scale. However, either only the longitudinal mode becomes strongly coupled, or the remaining modes of the theory become strongly coupled as well. From the perspective of self-interacting theory, the first case is natural. In cubic or quartic self-interactions, the transverse modes, dof which are also present in Maxwell’s electrodynamics, remain weakly coupled even after the perturbation theory for the longitudinal modes breaks down. The second outcome, in which the transverse modes, or tensor modes – modes of the metric perturbations – become strongly coupled, however, would be unphysical. It would indicate that in the presence of non-minimal coupling, one could either not be able to recover electromagnetism or discover gravitational waves.
By studying the Proca theory, non-minimally coupled to gravity in a flat background, and taking into account the metric perturbations around it, we will show that the last case takes place – both tensor and longitudinal modes become strongly coupled, yielding a physically inconsistent result. Moreover, we will show that the way to resolve this ambiguity lies in the introduction of the disformal coupling to the metric perturbations simultaneously. Such disformal transformations, which modify the original metric by adding a bilinear term in the external fields or their field-strength have gained a lot of attention in cosmology, and construction of new theories of gravity [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74]. In addition, we will show that the production of runaway modes – modes with negative frequency in curved background – recently reported in [26], can be cured with the same transformations.
@@ 2 The basics of linearized gravity and Proca theory
Here, the indices are raised and lowered with the Minkowski metric.
Let us now first study the degrees of freedom (dof) in gravity. For this, we will decompose the metric perturbations according to the spatial rotations:
@@ 3.1 $\beta$ coupling case
For the longitudinal modes, this means that the first term on the left-hand side of (28) is no longer vanishing. Once the scale $L_{\beta str}$ is reached, their first-order corrections should be evaluated as:
@@
Therefore, in this section, we have found that both tensor modes and longitudinal modes become strongly coupled once the scale reaches $L_{\beta str}$. However, this is an unphysical result. It would imply that no matter how small the mass of the photon is, gravitational waves would be strongly coupled if non-minimal coupling with the Ricci tensor is taken into account.
@@ 3.2 $\alpha$ coupling case
In the previous section we have seen that due to the $\beta$ interaction, both tensor and the longitudinal modes become strongly coupled. Let us now study the coupling that involves only the Ricci scalar. We can notice that at cubic order, only the gravitational potentials are included in the interaction:
@@
Therefore, in contrast to the previous case, we can safely avoid the danger of tensor modes becoming strongly coupled in this order. Thus, the only question to be answered is at which scale will the longitudinal modes enter a strong coupling regime. For simplicity, let us set the tensor and transverse vector modes to zero. Then, the most dominant terms in the Lagrangian are given by:
@@ 4 How to cure the tensor modes?
In the previous sections, we have studied the $\alpha$ and $\beta$ models, which involved the coupling of the vector field with the Ricci scalar and the Ricci tensor respectively. In the case of the $\alpha$ coupling, we have found that only the longitudinal modes become strongly coupled, due to the quartic self-interaction that appears after one resolves the constraints. In contrast, the coupling with the Ricci tensor proved to be troubling. There, we have found that both tensor and longitudinal modes become strongly coupled, which is a physically unacceptable result.
@@
Moreover, the couplings between $\phi$, $\psi$, and $S_{i}$ and the vector modes in the last line of $\mathcal{L}_{hA}$ can be neglected as well. One can show that the longitudinal mode in these relations enters in combination with the $A_{0}$ component, such that once the temporal constraint is resolved, it is multiplied by the mass squared, making such terms sub-dominant. Therefore, the most important terms are given by:
@@
for the tensor modes. Here, we have omitted the contribution of the longitudinal modes, which is now subdominant. By taking into account that the minimal amplitude of quantum fluctuations for the tensor modes is singular in the Planck mass, it is clear that the two types of modes will become strongly coupled only at the Planck scale, before which the perturbative approach still holds.
@@ 5 Resolving the Runaway Modes
in the lowest order, while the last term represents higher-order terms that involve couplings between the vector fields with the Ricci scalar or Ricci tensor, this guarantees that now $M_{S}^{2}=M_{T}^{2}$, and thus the runaway case disappears. It should be stressed that although not explicitly analyzed, the action (79), indicates that the theory in this disformal frame will have no strong coupling ambiguity in the tensor or runaway modes even for an arbitrary background.
@@ 6 Conclusion
In this work, we have studied Proca theory, non-minimally coupled to gravity. The cosmological aspects of this theory have recently drawn a lot of attention, making it an intriguing possibility to underline the gravitational production of dark matter and a possible candidate for generating primordial magnetic fields and inflationary candidates. By studying this theory in flat space-time, with metric perturbations also taken into account, we have found a somewhat surprising result. On the one hand, the interaction of the vector field with the Ricci scalar results in the strong coupling of the longitudinal modes. This is relatively natural to expect, as one obtains self-interaction of the longitudinal modes at the quartic order. Beyond it, the vector and tensor modes remain weakly coupled, up to the Planck length, due to the amplitude of quantum fluctuations for the tensor modes. On the other hand, the coupling of the vector modes with the Ricci tensor indicates that both tensor and longitudinal modes become strongly coupled at the same Vainsthein scale. This, however, is an unphysical result – it would imply that if one allows for such non-minimal coupling in nature, the gravitational waves would quickly lose their linear propagator. Moreover, a recent study has found a new class of the negative frequency modes in curved space-time, known as the runaway modes [26]. In this work, we have shown that its origin lies precisely in the coupling of the vector field to the Ricci tensor, which is problematic even in the flat background.
@@ Acknowledgements
I am grateful to Misao Sasaki for illuminating discussions and comments on the draft of the paper. I would also like to thank Edward "Rocky" Kolb for communicating his work on the Proca theory, and for very useful discussions. In addition, I would like to thank Elisa G. M. Ferreira, Viatcheslav Mukhanov, Shinji Mukohyama, Ryo Namba, and Ippei Obata for very useful discussions, and Evan McDonough for useful correspondence. The work of A. H. was supported by the World Premier International Research Center Initiative (WPI), MEXT, Japan.